OpenStudy (anonymous):

betsy, a recent retiree requires $6,000 per year in extra income. she has $50,000 to invest and can invest in b rated bonds paying 15% per year or in a certificate of deposit (CD) paying 7% per year. how much money should be invested in each to realize exactly $6,000 in interest per year? [please explain and show steps so i can study it ^_^]

OpenStudy (amistre64):

requires $6,000 per year has $50,000 to invest and can invest in b rated bonds paying 15% per year or in a certificate of deposit (CD) paying 7% per year. how much money should be invested in each to realize exactly $6,000 in interest per year? it a system of equations; set up b + c = 50000 15% b + 7% c = 6000 amount for b = (6000 - 50000(.07))/(.15-.07) maybe; its akin to the cramer rule

OpenStudy (amistre64):

If so, I get 31250 @ 15% bonds and the rest in cd